find x : 8x2+3x-6=0


1
Expert's answer
2021-05-20T08:10:02-0400

"F(x)=0"

"F(x)=8x^2+3x-6=0"


Now we can not find its root by normal factorization method;so we use here Shree Dharacharya Formula.

For this compare the equation "8x^2+3x-6" "=0" by "ax^2+bx+c=0"


From above comparison we have

"a=8,b=3,c=-6"


Hence root "="


"x=\dfrac{-b\pm \sqrt{b^2-4ac}}{2a}"


Putting the values of a,b,c in the above equation.


"x=\frac{-3\pm \sqrt{(-3)^2-4(8)(-6)}}{2\times8}"


"x=\dfrac{-3\pm\sqrt{9+192}}{16}"


"x=\dfrac{-3\pm\sqrt{201}}{16}"


"x=\dfrac{-3\pm13.85}{16}"


"x=\dfrac{-3+13.85}{16},\dfrac{-3-13.85}{16}"


"x=0.68,-1.05"


So the two roots of the equation are "0.68,-1.05" .